# Investigating Language Preference of Multilingual RAG Systems

### v2

@@ 1 Introduction
Beyond retrieval challenges and source conflicts, language preference is another critical issue in mRAG systems, often leading to inaccurate outputs. As illustrated in Case 1 of Figure 1, the retriever may prioritize particular languages—especially high-resource or query-language documents—at the expense of truly relevant information in low-resource language. Consequently, the Large Language Model (LLM) either produces an incorrect answer or deems the query unanswerable due to irrelevant content in the documents. Likewise, in Case 2, even if relevant documents are retrieved, the generator might favor passages in the query language or Latin scripts, ignoring essential evidence in lower-resource languages and resulting in inaccurate outputs. These preferences ultimately limit the effectiveness of mRAG, yielding biased rankings, reduced answer quality, and inconsistencies across languages Sharma et al. (2024).
@@
To address these questions, we present a comprehensive evaluation of language preferences throughout the entire mRAG pipeline across multiple languages. To systematically investigate the language preference problem of multilingual retrievers, we propose MultiLingualRank (MLR), a novel metric that quantifies language preference at the retriever level by measuring the shift in document rankings when non-query-language passages are translated into the query language. Our extensive experiments with diverse language combinations demonstrate that the retriever strongly prefers documents that are in high-resource languages and also share the same language as the query, confirming the presence of significant preference (§4).
@@
Figure 2: Overall framework of calculating MLR. For simplicity, we only consider three documents to calculate the MLR score.
2 MultiLingualRank
To evaluate the language preference of a multilingual retriever in the mRAG system, we introduce MultiLingualRank (MLR), a novel metric that quantifies how much the ranking of retrieved documents improves when non-query language documents are translated into the query language. As shown in Figure 2, MLR is computed in three stages: (i) retrieving documents across multiple languages, (ii) translating documents that are not in the query language into query language, and (iii) re-ranking the translated documents to measure rank improvements.
@@ 2.2 Stage 2: Translation of Non-Query Language Documents
2.3 Stage 3: Re-Ranking and MLR Score Computation
@@
The query-specific MLR score is then calculated as:
$\text{MLR}_{q}=\begin{cases}\displaystyle\frac{\Delta r_{q}}{\Delta r_{q}^{\max}}\times 100,&\text{if }\Delta r_{q}^{\max}>0,\\[6.0pt] 0,&\text{otherwise}.\end{cases}$
Finally, the overall MultiLingualRank is obtained by averaging the scores over all queries:
$\text{MLR}=\frac{1}{\lvert Q\rvert}\sum_{q\in Q}\text{MLR}_{q}.$
@@ 3.2 Models
Following previous work Chirkova et al. (2024), we use a strong and publicly available BGE-m3 Chen et al. (2024) as our multilingual retriever which can encode various languages we consider in our experiments. Consistent with the retriever, we use BGE-m3 Chen et al. (2024) as a re-ranking encoder for computing MLR. In addition, we use two other Sentence-BERT series re-ranking encoders Reimers and Gurevych (2019), paraphrase-multilingual-MiniLM-L12-v2 and paraphrase-multilingual-mpnet-base-v2 to generalize the experimental results. We abbreviate them as p-mMiniLM and p-mMpNet for better visibility of the table.
@@
We use recently released various strong multilingual LLM, aya-expanse-8b Dang et al. (2024) that can deal with various languages well. Also, we use strong LLMs, qwen 2.5-7B Instruct Team (2024) and Phi-4 14B Abdin et al. (2024), Llama-3.1-8B-Instruct Dubey et al. (2024) as our generators.
@@ 3.3 Other Implementation Details
Table 1: Language preference measured by MLR with different re-ranking encoders for various query–document language pairs. The $L_{q}=L_{d}$ column shows scores for matching query and document languages, while the remaining columns represent cross-lingual scenarios. Parentheses indicate the change from the $L_{q}=L_{d}$ column (positive for improvement, negative for decline). The highest score per row is in bold, and the second highest is underlined.
@@ 4.1.1 Experimental Setup
We evaluate eight language pairs under two scenarios: (1) monolingual settings where $L_{q}=L_{d}$, and (2) cross-lingual settings where $L_{q}\neq L_{d}$. In each case, our primary metric is MLR (MultiLingualRank), computed using three re-ranker encoders (bge-m3, p-mMiniLM, and p-mMpNet). For example, if the query is in English (en) but the target translation is in Korean (ko), we translate all non-English passages into Korean and then measure the rank changes with MLR.
@@ 4.1.3 Results for Cross-Lingual Settings ($L_{q}\neq L_{d}$)
Lower Overall MLR in Cross-Lingual Matching.
When $L_{q}\neq L_{d}$, the retriever performs cross-lingual matching, which typically results in lower MLR values than in monolingual cases. As indicated by the right-hand columns in Table 1 (highlighted in blue), cross-lingual setups are generally less preferred than their monolingual counterparts—except in cases involving English.
@@ 4.2.1 Experimental Setup
We also investigate whether the volume of available language resources affects MLR. We categorize languages into three groups based on their distribution in the pre-training corpus of recent LLMs: high-resource (e.g., English), mid-resource (e.g., Spanish), and low-resource (e.g., Korean). We use the same query set across all setups while systematically varying $L_{q}$ and $L_{d}$.
@@ 4.2.2 Results
The resource level of the query language ($L_{q}$) has a limited effect on cross-lingual preference. As shown in Table 1, when $L_{q}$ is high-resource (e.g., English), strong preference is observed only if $L_{d}$ also matches a high-resource language. Otherwise, the MLR scores remain similar regardless of whether $L_{q}$ is high-, mid-, or low-resource.
@@
In contrast, the language resource level of the document language ($L_{d}$) has a pronounced impact on MLR. As shown in Table 1, documents from high-resource languages consistently achieve the highest preference scores, followed by mid-resource and then low-resource languages. This trend (High > Mid > Low) suggests that extensive pre-training on high-resource languages enables stronger alignment, yielding higher MLR even across diverse query languages. Conversely, low-resource datastores typically produce lower MLR scores unless the query language also corresponds to that low-resource setting.
@@ 5.1.1 Experimental Setup
To assess the generator’s language preference, we measure answer consistency across eight languages: English (en), Korean (ko), Chinese (zh), French (fr), Japanese (ja), Italian (it), Portuguese (pt), and Spanish (es). For a given query, the generator produces responses in each of these languages using the same retrieved document set from multilingual datastores. We then compute the embedding similarity between each pair of generated answers, resulting in an 8×8 similarity matrix. We define the preference for a specific language as the average similarity score of the responses in that language.
@@ 5.2.1 Experimental Setup
To examine how language preference impacts overall mRAG performance, we isolate language effects by providing generators with retrieved passages unified in a single target language—selected from the eight candidates (en, ko, zh, fr, ja, it, pt, es). We retrieve data from multilingual sources, enabling a direct comparison between language preference (measured by MLR, as shown in Table 1) and performance across three query languages (en, ko, zh).
We evaluate four generators (aya-expanse-8B, Phi-4, Qwen2.5-7B-Instruct, and Llama3.1-8B-Instruct) using character 3-gram recall Chirkova et al. (2024) under three configurations. In the first configuration, passages are retrieved from multilingual resources, denoted as all. In the second, all retrieved passages are unified into a single target language (single-language document set). Finally, in the third configuration, we employ our proposed DKM-RAG framework (detailed in Section 6) to mitigate language preference. We also compute the average MLR score (across different query languages) for each language to indicate its overall preference.
@@
MLR (Preference) - 47.70 35.90 35.47 37.94 37.59 37.66 37.15 37.97 -
Table 2: Performance comparison between DKM-RAG and single/all language retrieval settings, showing character 3-gram recall scores for three query languages ($L_{q}\in\{\text{en},\text{ko},\text{zh}\}$) and eight passage languages. The bottom row shows average preference (MLR) scores. We highlight the cells corresponding to matching query and passage languages with a yellow background. The highest score per row is in bold, and the second highest is underlined.
@@ 6 Dual Knowledge Multilingual RAG
Based on this insight, we propose Dual Knowledge Multilingual RAG (DKM-RAG), a framework that leverages both external translated passage and internal knowledge as shown in Figure 4. First (#1), we retrieve documents for a given query from the all strategy and re-rank them. Next (#2), we obtain external translated passages, $P_{\text{translated}}$ by translating into the query language. And (#3), the rewriter LLM refines each translated passage in the context of the given query to produce refined passages, $P_{\text{refined}}$. This refining process utilizes a prompt to guide the model in integrating its internal knowledge, removing redundancy, and highlighting relevant information in a coherent and consistent style. For detailed prompts, please refer to Appendix B. Finally (#4), We concatenate the two sets to form the final passage set as input to the generator LLM, ensuring that responses are both contextually enriched and linguistically aligned with the query:
$P_{\text{final}}=\text{concat}\bigl{(}P_{\text{translated}},P_{\text{refined}}\bigr{)}.$
@@ 7.1 Multilingual RAG
Researchers explore challenges in mRAG such as problem of cross-lingual dense passage retrieval for low-resource languages Wu et al. (2024a) and research various techniques to address key challenges in mRAG, such as enhancing the performance of language models in low-resource languages Deshpande et al. (2024), resolving low-resource scenarios Zhang et al. (2024), and adapting language models for multilingual reasoning tasks Yoon et al. (2024). Benchmarks like MMTEB Enevoldsen et al. (2025) enable systematic evaluation of multilingual retrieval.
@@ 8 Conclusion
MLR entails translation and re-ranking steps. While this approach offers a principled way to quantify language preference, it also adds latency and computational cost, especially when dealing with large-scale multilingual corpora or real-time systems.
@@ Acknowledgments
We would like to thank Byeongjeong Kim for his comments and feedback about our figures. We also thank Gyutae Park for his minor correction of this work. This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
@@ Appendix E Language Preference of Other Languages
We also perform additional experiments to explore language preferences for languages not covered in Table 1, using the MLR score that we propose. As shown in Table 9, similar to the results in Table 1, the highest preferences are typically observed when $L_{q}=L_{d}$ across all query languages. English is also the most preferred language. For clarity, we omit results for other languages.
For most languages, such as Arabic, Finnish, German, and Russian, switching to a cross-lingual setup leads to a significant drop in MLR. For example, Arabic queries using the bge-m3 encoder achieve a monolingual score of 40.39, but cross-lingual retrieval (e.g., with Thai) results in a 6.80-point decrease.
@@
Table 9: Language preference measured by MLR with various re-ranking encoders for various query and document language combinations in a multilingual retriever. The $L_{q}=L_{d}$ column reports the diagonal scores where the query language matches the translated document language, while the remaining columns represent cross-lingual scenarios (i.e., where the query language differs from the document language). Scores in parentheses indicate the difference from the diagonal value (positive for an improvement, negative for a decline). The highest score for each row is highlighted in bold, and the second highest is underlined.
@@ Appendix G Case study
MLR
We provide an example of a document that improved MLR score, where the rank of a relevant document significantly increases after translation. In Table 10, the user query "영국 캐리비안에 언제 노예제가 폐지됐나요? (When was slavery abolished in the British Caribbean?)" is in Korean, whereas the original passage is in English. Initially, the document’s rank ($\mathbf{r_{d}^{\text{init}}}=34$) was relatively low, but after translating the passage into Korean and re-ranking ($\mathbf{r_{d}^{\text{re-rank}}}=2$), the document moved much closer to the top. This demonstrates how cross-lingual alignment can substantially improve retrieval performance in a multilingual setting. Notably, even if the passage content is semantically the same, language preference in the model can lead to poor alignment when the query and document are in different languages, adversely affecting retrieval. Translating the document into the query language effectively mitigates this issue.
@@
Table 10: An example of an improved MLR case. After translating the document into Korean, its rank improved from 34 to 2, illustrating language preference of retriever.

### v3

@@ 1 Introduction
Beyond retrieval challenges and source conflicts, language preference is another critical issue in mRAG systems, often leading to inaccurate outputs. As illustrated in Case 1 of Figure 1, the retriever may prioritize particular languages—especially high-resource or query-language documents—at the expense of truly relevant information in low-resource languages. Consequently, the Large Language Model (LLM) either produces an incorrect answer or deems the query unanswerable due to irrelevant content in the documents. Likewise, in Case 2, even if relevant documents are retrieved, the generator might favor passages in the query language or Latin scripts, ignoring essential evidence in lower-resource languages and resulting in inaccurate outputs. These preferences ultimately limit the effectiveness of mRAG, yielding biased rankings, reduced answer quality, and inconsistencies across languages Sharma et al. (2024).
@@
To address these questions, we present a comprehensive evaluation of language preferences throughout the entire mRAG pipeline across multiple languages. To systematically investigate the language preference problem of multilingual retrievers, we propose MultiLingualRankShift (MLRS), a novel metric that quantifies language preference at the retriever level by measuring the shift in document rankings when non-query-language passages are translated into the query language. Our extensive experiments with diverse language combinations demonstrate that the retriever strongly prefers documents that are in high-resource languages and also share the same language as the query, confirming the presence of significant preference (§4).
@@
Figure 2: Overall framework of calculating MLRS. For simplicity, we only consider three documents to calculate the MLRS score.
2 MultiLingualRankShift
To evaluate the language preference of a multilingual retriever in the mRAG system, we introduce MultiLingualRankShift (MLRS), a novel metric that quantifies how much the ranking of retrieved documents improves when non-query language documents are translated into the query language. As shown in Figure 2, MLRS is computed in three stages: (i) retrieving documents across multiple languages, (ii) translating documents that are not in the query language into the query language, and (iii) re-ranking the translated documents to measure rank improvements.
@@ 2.2 Stage 2: Translation of Non-Query Language Documents
2.3 Stage 3: Re-Ranking and MLRS Score Computation
@@
The query-specific MLRS score is then calculated as:
$\text{MLRS}_{q}=\begin{cases}\displaystyle\frac{\Delta r_{q}}{\Delta r_{q}^{\max}}\times 100,&\text{if }\Delta r_{q}^{\max}>0,\\[6.0pt] 0,&\text{otherwise}.\end{cases}$
Finally, the overall MultiLingualRankShift is obtained by averaging the scores over all queries:
$\text{MLRS}=\frac{1}{\lvert Q\rvert}\sum_{q\in Q}\text{MLRS}_{q}.$
@@ 3.2 Models
Following previous work Chirkova et al. (2024), we use a strong and publicly available BGE-m3 Chen et al. (2024) as our multilingual retriever which can encode various languages we consider in our experiments. Consistent with the retriever, we use BGE-m3 Chen et al. (2024) as a re-ranking encoder for computing MLRS. In addition, we use two other Sentence-BERT series re-ranking encoders Reimers and Gurevych (2019), paraphrase-multilingual-MiniLM-L12-v2 and paraphrase-multilingual-mpnet-base-v2 to generalize the experimental results. We abbreviate them as p-mMiniLM and p-mMpNet for better visibility of the table.
@@
We use a recently released strong multilingual LLM, aya-expanse-8b Dang et al. (2024) that can deal with various languages well. Also, we use strong LLMs, qwen 2.5-7B Instruct Team (2024) and Phi-4 14B Abdin et al. (2024), Llama-3.1-8B-Instruct Dubey et al. (2024) as our generators.
@@ 3.3 Other Implementation Details
Table 1: Language preference measured by MLRS with different re-ranking encoders for various query–document language pairs. The $L_{q}=L_{d}$ column shows scores for matching query and document languages, while the remaining columns represent cross-lingual scenarios. Parentheses indicate the change from the $L_{q}=L_{d}$ column (positive for improvement, negative for decline). The highest score per row is in bold, and the second highest is underlined.
@@ 4.1.1 Experimental Setup
We evaluate eight language pairs under two scenarios: (1) monolingual settings where $L_{q}=L_{d}$, and (2) cross-lingual settings where $L_{q}\neq L_{d}$. In each case, our primary metric is MLRS (MultiLingualRankShift), computed using three re-ranker encoders (bge-m3, p-mMiniLM, and p-mMpNet). For example, if the query is in English (en) but the target translation is in Korean (ko), we translate all non-English passages into Korean and then measure the rank changes with MLRS.
@@ 4.1.3 Results for Cross-Lingual Settings ($L_{q}\neq L_{d}$)
Lower Overall MLRS in Cross-Lingual Matching.
When $L_{q}\neq L_{d}$, the retriever performs cross-lingual matching, which typically results in lower MLRS values than in monolingual cases. As indicated by the right-hand columns in Table 1 (highlighted in blue), cross-lingual setups are generally less preferred than their monolingual counterparts—except in cases involving English.
@@ 4.2.1 Experimental Setup
We also investigate whether the volume of available language resources affects MLRS. We categorize languages into three groups based on their distribution in the pre-training corpus of recent LLMs: high-resource (e.g., English), mid-resource (e.g., Spanish), and low-resource (e.g., Korean). We use the same query set across all setups while systematically varying $L_{q}$ and $L_{d}$.
@@ 4.2.2 Results
The resource level of the query language ($L_{q}$) has a limited effect on cross-lingual preference. As shown in Table 1, when $L_{q}$ is high-resource (e.g., English), strong preference is observed only if $L_{d}$ also matches a high-resource language. Otherwise, the MLRS scores remain similar regardless of whether $L_{q}$ is high-, mid-, or low-resource.
@@
In contrast, the language resource level of the document language ($L_{d}$) has a pronounced impact on MLRS. As shown in Table 1, documents from high-resource languages consistently achieve the highest preference scores, followed by mid-resource and then low-resource languages. This trend (High > Mid > Low) suggests that extensive pre-training on high-resource languages enables stronger alignment, yielding higher MLRS even across diverse query languages. Conversely, low-resource datastores typically produce lower MLRS scores unless the query language also corresponds to that low-resource setting.
@@
4.3 Impact of Translation Quality
To investigate the effect of translation quality on language preference as measured by the MLRS metric, we conduct a small-scale experiment by randomly sampling 100 queries out of the full set of 2,827. We translate these queries using GPT-4o-mini, which demonstrated highly robust translation quality. We evaluate five languages (English, Korean, Spanish, Chinese, and French) and re-ranked the top-10 retrieved documents using the BGE-M3 encoder, and compute rank shifts within the top-10 to derive MLRS for monolingual ($L_{q}=L_{d}$) and cross-lingual ($L_{q}\neq L_{d}$) settings. The results, shown in Table 2, are consistent with prior MLRS findings and further reveal that GPT-based translation amplifies the distinctness of language preference trends, as indicated by larger inter-language preference gaps.
$L_{q}=L_{d}$ $L_{q}\neq L_{d}$
Query Lang. en ko zh fr es
en 58.29 34.98 (-23.31) 36.10 (-22.19) 38.49 (-19.80) 40.07 (-18.22)
ko 47.15 42.59 (-4.56) 37.50 (-9.65) 39.31 (-7.84) 38.74 (-8.41)
zh 47.42 46.66 (-0.76) 36.89 (-10.53) 38.24 (-9.18) 38.10 (-9.32)
fr 47.35 50.76 (+3.41) 34.90 (-12.45) 36.10 (-11.25) 38.90 (-8.45)
es 45.98 50.73 (+4.75) 35.06 (-10.92) 36.18 (-9.80) 40.52 (-5.46)
Table 2: MLRS scores for monolingual ($L_{q}=L_{d}$) and cross-lingual ($L_{q}\neq L_{d}$) settings on 100 randomly sampled queries.
@@ 5.1.1 Experimental Setup
To assess the generator’s language preference, we measure multilingual answer consistency across eight languages: English (en), Korean (ko), Chinese (zh), French (fr), Japanese (ja), Italian (it), Portuguese (pt), and Spanish (es). As shown in Table 13, for a given query the generator produces responses in each language using the same set of retrieved documents from the multilingual datastore. We then compute the embedding similarity between each pair of generated answers, resulting in an 8×8 similarity matrix. We define the preference for a specific language as the average similarity score of the responses in that language.
@@ 5.2.1 Experimental Setup
To examine how language preference impacts overall mRAG performance, we isolate language effects by providing generators with retrieved passages unified in a single target language—selected from the eight candidates (en, ko, zh, fr, ja, it, pt, es). We retrieve data from multilingual sources, enabling a direct comparison between language preference (measured by MLRS, as shown in Table 1) and performance across three query languages (en, ko, zh).
We evaluate four generators (aya-expanse-8B, Phi-4, Qwen2.5-7B-Instruct, and Llama3.1-8B-Instruct) using character 3-gram recall Chirkova et al. (2024) under three configurations. In the first configuration, passages are retrieved from multilingual resources, denoted as all. In the second, all retrieved passages are unified into a single target language (single-language document set). Finally, in the third configuration, we employ our proposed DKM-RAG framework (detailed in Section 6) to mitigate language preference. We also compute the average MLRS score (across different query languages) for each language to indicate its overall preference.
@@
MLRS (Preference) - 47.70 35.90 35.47 37.94 37.59 37.66 37.15 37.97 -
Table 3: Performance comparison between DKM-RAG and single/all language retrieval settings, showing character 3-gram recall scores for three query languages ($L_{q}\in\{\text{en},\text{ko},\text{zh}\}$) and eight passage languages. The bottom row shows average preference (MLRS) scores. We highlight the cells corresponding to matching query and passage languages with a yellow background. The highest score per row is in bold, and the second highest is underlined.
@@ 6 Dual Knowledge Multilingual RAG
Based on this insight, we propose Dual Knowledge Multilingual RAG (DKM-RAG), a framework that leverages both external translated passages and internal knowledge as shown in Figure 4. First (#1), we retrieve documents for a given query from the all strategy and re-rank them. Next (#2), we obtain external translated passages, $P_{\text{translated}}$ by translating into the query language. And (#3), the rewriter LLM refines each translated passage in the context of the given query to produce refined passages, $P_{\text{refined}}$. This refining process utilizes a prompt to guide the model in integrating its internal knowledge, removing redundancy, and highlighting relevant information in a coherent and consistent style. For detailed prompts, please refer to Appendix B. Finally (#4), We concatenate the two sets to form the final passage set as input to the generator LLM, ensuring that responses are both contextually enriched and linguistically aligned with the query.
@@ Results.
Ablation Study.
To prove the effectiveness of concatenating translated passages and refined passages in the DKM-RAG framework, we provide an ablation study of each component in DKM-RAG. As stated in Table 4, removing any component from DKM-RAG decreases performance, highlighting that every part is crucial to its effectiveness.
@@ 7.1 Multilingual RAG
Researchers explore challenges in mRAG, such as the problem of cross-lingual dense passage retrieval for low-resource languages Wu et al. (2024a), and propose various techniques to address key challenges in mRAG, such as enhancing the performance of language models in low-resource languages Deshpande et al. (2024), resolving low-resource scenarios Zhang et al. (2024), and adapting language models for multilingual reasoning tasks Yoon et al. (2024). Benchmarks like MMTEB Enevoldsen et al. (2025) enable systematic evaluation of multilingual retrieval.
@@ 7.2 Language Preference in mRAG
DKM-RAG w/o $P_{\text{refined}}$ w/o $P_{\text{translated}}$
$L_{q}=$ en
aya-expanse-8b 82.60 79.34 81.10
Phi-4 82.59 78.89 81.08
Qwen2.5-7B-Instruct 82.60 79.11 81.06
Llama3.1-8B-Instruct 82.57 79.28 81.19
$L_{q}=$ zh
aya-expanse-8b 44.57 38.31 39.44
Phi-4 44.56 36.76 38.95
Qwen2.5-7B-Instruct 44.70 38.31 39.78
Llama3.1-8B-Instruct 44.51 38.48 39.35
$L_{q}=$ ko
aya-expanse-8b 55.01 49.66 46.15
Phi-4 54.82 49.25 45.24
Qwen2.5-7B-Instruct 54.85 49.44 45.32
Llama3.1-8B-Instruct 54.99 49.87 45.55
@@ 8 Conclusion
MLRS entails translation and re-ranking steps. While this approach offers a principled way to quantify language preference, it also adds latency and computational cost, especially when dealing with large-scale multilingual corpora or real-time systems.
@@ Acknowledgments
We would like to thank Byeongjeong Kim for his comments and feedback about our figures. We also thank Gyutae Park for his minor corrections to this work. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)].
@@ Appendix E Language Preference of Other Languages
We also perform additional experiments to explore language preferences for languages not covered in Table 1, using the MLRS score that we propose. As shown in Table 11, similar to the results in Table 1, the highest preferences are typically observed when $L_{q}=L_{d}$ across all query languages. English is also the most preferred language. For clarity, we omit results for other languages.
For most languages, such as Arabic, Finnish, German, and Russian, switching to a cross-lingual setup leads to a significant drop in MLRS. For example, Arabic queries using the bge-m3 encoder achieve a monolingual score of 40.39, but cross-lingual retrieval (e.g., with Thai) results in a 6.80-point decrease.
@@
Table 11: Language preference measured by MLRS with various re-ranking encoders for various query and document language combinations in a multilingual retriever. The $L_{q}=L_{d}$ column reports the diagonal scores where the query language matches the translated document language, while the remaining columns represent cross-lingual scenarios (i.e., where the query language differs from the document language). Scores in parentheses indicate the difference from the diagonal value (positive for an improvement, negative for a decline). The highest score for each row is highlighted in bold, and the second highest is underlined.
@@ Appendix G Case study
MLRS
We provide an example of a document that improved MLRS score, where the rank of a relevant document significantly increases after translation. In Table 12, the user query "영국 캐리비안에 언제 노예제가 폐지됐나요? (When was slavery abolished in the British Caribbean?)" is in Korean, whereas the original passage is in English. Initially, the document’s rank ($\mathbf{r_{d}^{\text{init}}}=34$) was relatively low, but after translating the passage into Korean and re-ranking ($\mathbf{r_{d}^{\text{re-rank}}}=2$), the document moved much closer to the top. This demonstrates how cross-lingual alignment can substantially improve retrieval performance in a multilingual setting. Notably, even if the passage content is semantically the same, language preference in the model can lead to poor alignment when the query and document are in different languages, adversely affecting retrieval. Translating the document into the query language effectively mitigates this issue.
@@
Table 12: An example of an improved MLRS case. After translating the document into Korean, its rank improved from 34 to 2, illustrating language preference of retriever.